Suppose that median disposable income measured
in current dollars is $10,000 and $20,000 for 1978
and 1988, respectively. The economy's price
index is 100 in the base year of 1978 and 250 in
1988. Which of the following is true?
(A) Inflation, as measured by the price index, rose
250 percent over the period.
(B) Nominal median disposable income rose 50
percent over the period.
(C) Real median disposable income rose 250
percent over the period.
(D) Real median disposable income fell 50 percent
over the period.
(E) The $10,000 in 1978, measured in 1988
dollars, is equivalent to $25,000
1
Expert's answer
2020-07-28T15:02:20-0400

The correct answer is E.

This is calculated as follows:


"\frac{CPI_{1988}}{CPI_{1978}} \times 1978 Value"


"\frac{250}{100}\times 10,000 = 25,000"